# Program evaluation with remotely sensed outcomes

### v2

@@ 3.3 Main result: A novel formula
Our main result nonparametrically identifies the ATE in the experimental sample, without a parametric model that restricts the distribution or dimensionality of the RSV. We derive a novel formula for combining experimental and observational data.
@@
To begin, we express the RSV distribution in the experimental sample as a mixture of potential outcomes. The mixture weights are identified by the observational sample.
Lemma 3.1 (Identification as generative model).
@@
Suppose Assumptions 2.12.2, and 2.3(ii) hold. Then, for any $d\in\{0,1\},$ $x\in\mathcal{X}$, and $r\in\mathcal{R}$, $\delta^{e}_{d}(r,x)=\left\{\delta_{1}^{o}(r,x)-\delta_{0}^{o}(r,x)\right\}\mu(d,x)+\delta^{o}_{0}(r,x)$ where $\mu(d,x):=\mathbb{E}\{Y(d)\mid S=e,X=x\}$ is the conditional average potential outcome in the experiment, while $\delta_{d}^{e}(r,x):=\Pr(R=r\mid S=e,X=x,D=d)$ and $\delta_{y}^{o}(r,x):=\Pr(R=r\mid S=o,X=x,Y=y)$ are RSV distributions.
@@
By Lemma 3.1, we recover the ATE in the experimental sample by combining (i) how the RSV varies with the treatment in the experimental sample, with (ii) how the RSV varies with the outcome in the observational sample. This combination leverages our key assumption: stability of the RSV across samples.
@@
While Lemma 3.1 identifies the ATE in the experimental sample, it suggests a challenging estimation problem: it involves the conditional distribution of the high dimensional RSV. One possible way forward is to develop a complex parametric model, e.g. a generative model of the satellite image distribution conditional upon poverty measurements or environmental outcomes. Even if such a generative model could be developed, it would be prone to mis-specification.
@@
We follow a different path that does not require a generative model on $R$. We use Bayes’ rule to rewrite Lemma 3.1 as a conditional moment equation. This transformation avoids estimation of the RSV’s conditional distribution and recovers a classic econometric estimation problem.
@@
Theorem 3.1 (Identification as conditional moment).
Under the conditions of Lemma 3.1,
[…]
$\Delta^{e}(x):=\frac{1\{D=1,S=e\}}{\Pr(D=1,S=e\mid X=x)}-\frac{1\{D=0,S=e\}}{\Pr(D=0,S=e\mid X=x)}$, $\Delta^{o}(x):=\frac{1\{Y=1,S=o\}}{\Pr(Y=1,S=o\mid X=x)}-\frac{1\{Y=0,S=o\}}{\Pr(Y=0,S=o\mid X=x)}$.
@@
Theorem 3.1 identifies the treatment effect as the solution to a set of conditional moment equalities. Intuitively, the conditional average treatment effect $\theta(x)$ balances treatment variation from the experimental sample $\Delta^{e}(x)$ and outcome variation from the observational sample $\Delta^{o}(x)$, so that their projections onto the remotely sensed variable $R$ match. Consequently, we can leverage a celebrated literature on conditional moment equalities e.g. Chamberlain, (1987); Newey, (1993) for estimation and inference.
@@
For identification, Theorem 3.1 implies that we may introduce representations of the RSV. Such representations can be arbitrary, as long as they predict outcome variation.
Corollary 3.1 (Identification as representation).
@@
Under Lemma 3.1’s conditions, $\theta(x)=\frac{\mathbb{E}\{H(x,R)\Delta^{e}(x)\mid X=x\}}{\mathbb{E}\{H(x,R)\Delta^{o}(x)\mid X=x\}}$ for any representation $H(x,R)$ with $\mathbb{E}\{H(x,R)\Delta^{o}(x)\mid X=x\}\neq 0$.
By Corollary 3.1, many representations of the RSV provide identification. However, naive choices may be inefficient, producing needlessly large standard errors. Section 4 asks: what is the optimal representation of the RSV for downstream causal inference? Our answer develops a connection between “representation learning” e.g. Johannemann et al., (2019); Vafa et al., (2024) and classical results for conditional moment equalities.
Remark 3.2 (Testable implication).
@@
Our main identifying assumptions are jointly testable. By Corollary 3.1, any predictive representation $H(x,R)$ identifies $\theta(x)$. If different representations yield significantly different estimates, then we can reject our identifying assumptions. Future work may develop bounds under violations of our identifying assumptions.
@@
So far, we have focused on the setting of Assumption 2.3(ii), which allows incomplete cases but disallows direct effects of the treatment on the RSV. Our main result also holds for the setting of Assumption 2.3(i), which requires complete cases yet allows direct effects of the treatment on the RSV. Recall that $\mu(d,x):=\mathbb{E}\{Y(d)\mid S=e,X=x\}$ is the conditional average potential outcome in the experimental sample.
Theorem 3.2 (Identification as conditional moment with direct effects).
Suppose Assumptions 2.1, 2.2, and 2.3(i) hold. Then, for any $d\in\{0,1\}$ and any $x\in\mathcal{X}$,
[…]
@@
Once again, the causal estimand reconciles treatment variation from the experimental sample with outcome variation from the observational sample, so that their projections onto the remotely sensed variable $R$ match. Theorem 3.2 allows direct effects; the key assumption in our framework is stability in Assumption 2.2. Corollary 3.1 and Remark 3.2 extend accordingly.
Remark 3.3 (Some experimental outcomes).
@@
Remark 3.4 (Multi-valued outcomes).
Our identification, estimation, and inference results generalize to discrete and continuous outcomes. Appendix E extends our identification result to discrete outcomes. We generalize the conditional moment equations. Estimation and inference remain essentially the same, under a minimum rank condition that requires the RSV to predict each outcome value well.
Appendix F extends our results to continuous outcomes. We describe how researchers can conduct estimation and inference using a discrete approximation, and characterize the worst-case bias of that discrete approximation. Intuitively, more complexity of $\mathcal{Y}$ must be compensated by additional regularity elsewhere.
@@ 4 Estimation and inference
As a secondary contribution, we demonstrate that our main contribution (Theorem 3.1) implies an optimal representation of the RSV for downstream causal estimation and inference. The techniques are standard; we simply point out and interpret the connection between modern remote sensing and classical conditional moment analysis. Using well known techniques, we derive valid $n^{-1/2}$ inference without rate conditions and without complexity restrictions on the researcher’s RSV-based predictions, justifying the use of complex deep learning algorithms.
@@
For clarity, we focus on the case in which the outcome is binary, there are no pretreatment covariates, and Assumption 2.3(ii) holds. A more general case is straightforward to implement by incorporating additional moment conditions, as discussed in Appendix E.
4.1 Optimal representation for program evaluation
In this setting, Theorem 3.1 and Corollary 3.1 simplify. The causal parameter satisfies $\mathbb{E}(\Delta^{e}-\Delta^{o}\theta\mid R)=0$ and hence $\mathbb{E}\{(\Delta^{e}-\Delta^{o}\theta)H(R)\}=0$ for any predictive representation $H(R)$. The treatment and outcome variation simplify to scalars:
$\Delta^{e}=\frac{1\{D=1,S=e\}}{\Pr(D=1,S=e)}-\frac{1\{D=0,S=e\}}{\Pr(D=0,S=e)},\;\Delta^{o}=\frac{1\{Y=1,S=o\}}{\Pr(Y=1,S=o)}-\frac{1\{Y=0,S=o\}}{\Pr(Y=0,S=o)}.$ (1)
Since $(S,D,Y)$ are binary, the denominators can be estimated by simple counts.
While any predictive representation $H$ gives inference, different choices have different efficiency properties. We appeal to classical ideas in econometrics to derive the optimal representation $H^{*}$, which achieves the semiparametric efficiency bound. Write the conditional moment as the regression $\Delta^{e}=\Delta^{o}\theta+\epsilon$ with $\mathbb{E}(\epsilon\mid R)=0$. The optimal representation is then $H^{*}(R)=\frac{\mathbb{E}(\Delta^{o}\mid R)}{\sigma^{2}(\theta,R)}$, where $\sigma^{2}(\theta,R):=\mathbb{E}\{(\Delta^{e}-\Delta^{o}\theta)^{2}\mid R\}$ (Chamberlain,, 1987; Newey,, 1993).
For downstream causal inference, the optimal representation of the high dimensional remotely sensed variable $R\in\mathcal{R}$ is a scalar $H^{*}(R)\in\mathbb{R}$. Concretely, $H^{*}(R)$ is the optimal compression of a satellite image for the task of program evaluation. Interestingly, different causal estimands imply different optimal compressions.
@@ 4.1 Optimal representation for program evaluation
This connection has another consequence for empirical practice: to optimally use the RSV, we should not only predict the outcome $Y$ using observational data, but also predict the treatment $D$ using experimental data, and predict the sample indicator $S$ using all data. The first prediction is part of common practice, but the second and third are not. By using all three predictions, we use the RSV most efficiently.
@@
We confirm this connection by further developing the expression for $H^{*}(R)$: the numerator $\mathbb{E}(\Delta^{o}\mid R)$ contains $\Pr(Y=1,S=o|R)$ while the denominator $\sigma^{2}(\theta,R)$ contains $\Pr(D=1,S=e|R)$; see Lemma 4.1 below. Finally, by the definition of conditional probability, $\Pr(Y=1,S=o|R)=\Pr(Y=1|S=o,R)\Pr(S=o|R)$ and $\Pr(D=1,S=e|R)=\Pr(D=1|S=e,R)\Pr(S=e|R)$.
@@ 4.2 Inference with learned representations
For simplicity, we describe our inferential procedure using sample splitting with train and test folds, though our proofs in Appendix D allow for cross fitting with any fixed number of folds. We first state our inferential procedure at a high level before filling in the details:
@@
    Learn the optimal representation on train: $\widehat{H}(R)$.
@@
    • Construct an initial causal estimate, by regressing $\widehat{\mathbb{E}}(\Delta^{e}|R)$ on $\widehat{\mathbb{E}}(\Delta^{o}|R)$: $\widehat{\theta}_{\textsc{init}}$.
    • Combine these into the optimal representation: $\widehat{H}(R)$.
@@
    Construct an efficient causal estimate on test: $\widehat{\theta}$.
The overall structure is familiar (Angrist et al.,, 1999; Chernozhukov et al.,, 2018). To simplify the algorithm statement, we introduce some notation for simple events and an algebraic identity.
Let $\mathbb{E}_{\textsc{train}}=\frac{1}{|\textsc{train}|}\sum_{i\in\textsc{train}}(\cdot)$ and $\mathbb{E}_{\textsc{test}}=\frac{1}{|\textsc{test}|}\sum_{i\in\textsc{train}}(\cdot)$. Slightly abusing notation, the counting operation, $\textsc{count}_{\textsc{event}}$, can be $\mathbb{E}_{\textsc{train}}(1_{\textsc{event}})$ in the second step or $\mathbb{E}_{\textsc{test}}(1_{\textsc{event}})$ in the third step, which will be clear from the context.
Lemma 4.1.
$(\Delta^{e}-\Delta^{o}\theta)^{2}=\frac{1\{D=1,S=e\}}{\Pr(D=1,S=e)^{2}}+\frac{1\{D=0,S=e\}}{\Pr(D=0,S=e)^{2}}+\theta^{2}\left[\frac{1\{Y=1,S=o\}}{\Pr(D=Y,S=o)^{2}}+\frac{1\{Y=0,S=o\}}{\Pr(Y=0,S=o)^{2}}\right].$
Algorithm 4.1 (Inference).
@@
    Learn the optimal representation on train: $\widehat{H}(R)$.
@@
      Learn the optimal representation: $\widehat{H}(R)=\frac{\widehat{\mathbb{E}}(\Delta^{o}|R)}{\widehat{\sigma}^{2}(\widehat{\theta}_{\textsc{init}},R)}$ where $\widehat{\sigma}^{2}(\widehat{\theta}_{\textsc{init}},R)$ is constructed from the marginal probabilities, predictors, and initial estimate via Lemma 4.1.
@@
    Construct an efficient causal estimate on test: $\widehat{\theta}$.
@@
      Construct an efficient causal estimate: $\widehat{\theta}=\frac{\mathbb{E}_{\textsc{test}}\{\widehat{\Delta}^{e}\widehat{H}(R)\}}{\mathbb{E}_{\textsc{test}}\{\widehat{\Delta}^{o}\widehat{H}(R)\}}$ where $\widehat{\Delta}^{e}$ and $\widehat{\Delta}^{o}$ are constructed from marginal probabilities according to (1).
@@
      Bootstrap its confidence interval: $\widehat{\theta}\pm c_{\alpha}\hat{v}n^{-1/2}$, where $c_{\alpha}$ is the $21-\alpha/$ 2 quantile of the standard Gaussian and $\hat{v}n^{-1/2}$ is the bootstrap standard error of $\widehat{\theta}$ while fixing $\widehat{H}(R)$.
@@
Sample splitting may be eliminated under complexity restrictions that tolerate simple machine learning procedures. See e.g. Chernozhukov et al., (2020) for a recent summary.
4.3 Robustness to mis-specification
In empirical research, the RSV is typically unstructured and high dimensional, e.g. a satellite image. The prediction is typically conducted by a complex machine learning algorithm, e.g. a deep convolutional neural network. For this realistic setting, rates of convergence are often unknown. In other words, we have no reason to believe that $\textsc{pred}_{Y}(R)$ converges to $\Pr(Y=1|S=o,R)$, nor that $\widehat{H}(R)$ converges to $H^{*}(R)$. Even carefully crafted architectures positing a generative model would typically be mis-specified.
For this reason, we place a weaker regularity condition: the predictions, and hence the representation estimator, have some probability limit; they may be mis-specified.
Assumption 4.1 (Limit).

### v3

@@ 3.3 Main Result: An Identification Formula
Our main result nonparametrically identifies the ATE in the experimental sample without a parametric model that restricts the distribution or dimensionality of the RSV. We derive a novel formula for combining the experimental and observational samples.
@@
To begin, we express the RSV distribution in the experimental sample as an affine transformation of the potential outcomes we wish to identify. The slope and intercept depend on how the RSV relates to the outcome, and they are identified by the observational sample.
Lemma 1 (Identification as generative model).
@@
Suppose Assumptions 12, and 3(ii) hold. Then, for any $d\in\{0,1\},$ $x\in\mathcal{X}$, and $r\in\mathcal{R}$, $\delta^{e}_{d}(r,x)=\left\{\delta_{1}^{o}(r,x)-\delta_{0}^{o}(r,x)\right\}\mu(d,x)+\delta^{o}_{0}(r,x),$ where $\mu(d,x):=\mathbb{E}\{Y(d)\mid S=e,X=x\}$ is the conditional average potential outcome in the experimental sample. Here, $\delta_{d}^{e}(r,x):=f_{R}(r\mid S=e,X=x,D=d)$ and $\delta_{y}^{o}(r,x):=f_{R}(r\mid S=o,X=x,Y=y)$ are RSV conditional densities in the experimental and observational samples, respectively.
@@
By Lemma 1, we recover the ATE in the experimental sample by combining (i) how the RSV varies with the treatment in the experimental sample with (ii) how the RSV varies with the outcome in the observational sample. This combination leverages our key assumption: stability of the RSV across samples.
@@
While Lemma 1 identifies the ATE in the experimental sample, it suggests a challenging estimation problem: it involves the conditional distribution of the high-dimensional RSV. One path forward is to develop a complex parametric model, e.g., a generative model of the satellite image distribution conditional upon poverty measurements or environmental outcomes. Even if such a generative model could be developed, it would be prone to misspecification.
@@
We follow a different path that does not require a generative model for $R$. We use Bayes’ rule to rewrite Lemma 1 as a conditional moment equation. This transformation avoids estimation of the RSV’s conditional distribution and recovers a classic econometric estimation problem.
@@
Theorem 1 (Identification as conditional moment).
Under the conditions of Lemma 1,
$\text{for any $x\in\mathcal{X}$,}\quad\mathbb{E}\{\Delta^{e}(x)-\Delta^{o}(x)\theta(x)\mid X=x,R\}=0\quad\text{almost surely, where}$
$\Delta^{e}(x):=\frac{1\{D=1,S=e\}}{\Pr(D=1,S=e\mid X=x)}-\frac{1\{D=0,S=e\}}{\Pr(D=0,S=e\mid X=x)}$ and $\Delta^{o}(x):=\frac{1\{Y=1,S=o\}}{\Pr(Y=1,S=o\mid X=x)}-\frac{1\{Y=0,S=o\}}{\Pr(Y=0,S=o\mid X=x)}$.
@@
Theorem 1 identifies the treatment effect as the solution to a set of conditional moment equalities. The conditional average treatment effect $\theta(x)$ balances treatment variation from the experimental sample $\Delta^{e}(x)$ and outcome variation from the observational sample $\Delta^{o}(x)$, so that their projections onto the remotely sensed variable $R$ match. Consequently, we can leverage a celebrated literature on conditional moment equalities (e.g., Chamberlain, 1987; Newey, 1993) for estimation and inference.
For intuition on Theorem 1, consider the case without covariates. Then $\Delta^{e}$ is simply a scalar computed from the experimental sample, where the treatment is observed. Similarly, $\Delta^{o}$ is simply a scalar computed from the observational sample, where the outcome is observed. Theorem 1 shows that $\theta_{0}=\frac{\mathbb{E}(\Delta^{e}|R)}{\mathbb{E}(\Delta^{o}|R)}.$ This ratio appears to be a new formula for data combination, where the numerator and denominator are from different samples. The numerator reflects how both the treatment and the outcome affect the RSV, so we must divide it by the effect of the outcome on the RSV, in order to isolate the desired causal parameter.
@@
For identification, Theorem 1 implies that we may introduce representations of the RSV. Such representations can be arbitrary, as long as they predict outcome variation.
Corollary 1 (Identification as representation).
@@
Under Lemma 1’s conditions, $\theta(x)=\frac{\mathbb{E}\{H(x,R)\Delta^{e}(x)\mid X=x\}}{\mathbb{E}\{H(x,R)\Delta^{o}(x)\mid X=x\}}$ for any representation $H(x,R)$ with $\mathbb{E}\{H(x,R)\Delta^{o}(x)\mid X=x\}\neq 0$.
Remark 3 (Testable implication).
@@
Our main identifying assumptions are jointly testable. By Corollary 1, any predictive representation $H(x,R)$ identifies $\theta(x)$. If different representations yield significantly different estimates, then we can reject our identifying assumptions. Future work may develop bounds under violations of our identifying assumptions.
Once again, for intuition on Corollary 1, consider the scenario without covariates. In this case, $H(R)$ is simply a scalar representation of the RSV. Corollary 1 shows that $\theta_{0}=\frac{\mathbb{E}\{\Delta^{e}H(R)\}}{\mathbb{E}\{\Delta^{o}H(R)\}}$. Any representation $H(R)$ is valid as long as the representation is predictive: $\mathbb{E}\{\Delta^{o}H(R)\}\neq 0$. A weak RSV test asks whether $\mathbb{E}\{\Delta^{o}H(R)\}\approx 0$, which can be tested using existing frameworks for weak instruments. In the same spirit, a joint test of our identifying assumptions asks whether two representations $H(R)$ and $H^{\prime}(R)$ give similar estimates. Dissimilar estimates would be evidence against our assumptions.
By Corollary 1, many representations of the RSV provide identification. However, naive choices may be inefficient, producing needlessly large standard errors. Section 4 asks: what representation of the RSV should be chosen for precise, downstream causal inference? Our answer develops a connection between “representation learning” (e.g., Johannemann et al., 2019; Vafa et al., 2025) and classical results for conditional moment equalities.
@@
So far, we have focused on the setting of Assumption 3(ii), which allows incomplete cases but disallows direct effects of the treatment on the RSV. Our main result also holds for the setting of Assumption 3(i), which requires complete cases yet allows direct effects of the treatment on the RSV. Recall that $\mu(d,x):=\mathbb{E}\{Y(d)\mid S=e,X=x\}$ is the conditional average potential outcome in the experimental sample.
Theorem 2 (Identification as conditional moment with direct effects).
Suppose Assumptions 1, 2, and 3(i) hold. Then, for any $d\in\{0,1\}$ and any $x\in\mathcal{X}$,
$\mathbb{E}\{\widetilde{\Delta}^{e}(d,x)-\widetilde{\Delta}^{o}(d,x)\mu(d,x)\mid R,X=x\}=0\quad\text{almost surely, where}$
@@
Once again, the causal estimand reconciles treatment variation from the experimental sample with outcome variation from the observational sample, so that their projections onto the remotely sensed variable $R$ match. Theorem 2 allows direct effects; the key assumption in our framework is stability in Assumption 2. Corollary 1 and Remark 3 extend accordingly.
Remark 4 (Some experimental outcomes).
@@
Remark 5 (Multi-valued outcomes).
Our identification, estimation, and inference results generalize to discrete and continuous outcomes. Appendix E extends our identification result to discrete outcomes. We generalize the conditional moment equations. Estimation and inference remain essentially the same, under a minimum rank condition that requires the RSV to predict each outcome value well. Appendix F extends our results to continuous outcomes.
@@ 4 Estimation and Inference
As a secondary contribution, we demonstrate that our identification result provides guidance on how to choose a representation of the RSV for valid, precise, and robust inference on the causal parameter. We point out and interpret the connection between modern remote sensing and classical conditional moment analysis, which allows us to use standard techniques. We derive valid $n^{-1/2}$ inference without rate conditions and without complexity restrictions on the researcher’s RSV-based predictions, justifying the use of complex deep learning algorithms that may be misspecified.
@@
For clarity, we focus on the simple case in which the outcome is binary, there are no pretreatment covariates, and Assumption 3(ii) holds. We discuss the more general case with discrete outcomes and discrete or continuous covariates in Appendix E.
4.1 Choice of Representation for Program Evaluation
Without covariates, our main identification result in Theorem 1 simplifies. The treatment and outcome variation are simply scalars:
$\Delta^{e}=\frac{1\{D=1,S=e\}}{\Pr(D=1,S=e)}-\frac{1\{D=0,S=e\}}{\Pr(D=0,S=e)},\quad\Delta^{o}=\frac{1\{Y=1,S=o\}}{\Pr(Y=1,S=o)}-\frac{1\{Y=0,S=o\}}{\Pr(Y=0,S=o)}.$ (1)
Since $(S,D,Y)$ are binary, the denominators can be estimated by simple counts. The treatment effect is identified as $\theta=\frac{\mathbb{E}(\Delta^{e}|R)}{\mathbb{E}(\Delta^{o}|R)}$. Therefore, $\theta=\frac{\mathbb{E}\{\Delta^{e}H(R)\}}{\mathbb{E}\{\Delta^{o}H(R)\}}$ for any representation $H(R)$ that is predictive of outcome variation in the sense that $\mathbb{E}\{\Delta^{o}H(R)\}\neq 0$. While any predictive representation can be used for inference, different choices have different efficiency properties. We now ask, which representation should researchers use in practice?
By rearranging our identifying formula, we highlight the connection to classical ideas in econometrics and thereby derive an answer. Specifically, the treatment effect $\theta$ may be viewed as the coefficient in the regression model
$\Delta^{e}=\theta\Delta^{o}+\epsilon,\quad\mathbb{E}(\epsilon|R)=0$ (2)
where the residual $\epsilon:=\Delta^{e}-\theta\Delta^{o}$ is mean zero conditional on $R$ by Theorem 1. Results in Chamberlain (1987) and Newey (1993) demonstrate that the representation that achieves efficiency within a class of models satisfying (2) is $H^{*}(R)=\frac{\mathbb{E}(\Delta^{o}\mid R)}{\sigma^{2}(\theta,R)}$, where $\sigma^{2}(\theta,R):=\mathbb{E}\{(\Delta^{e}-\Delta^{o}\theta)^{2}\mid R\}$.
For downstream causal inference, we find that the efficient representation of the high-dimensional remotely sensed variable $R\in\mathcal{R}$ is a simple, continuous scalar $H^{*}(R)\in\mathbb{R}$. The formula for $H^{*}(R)$ is a conditional expectation divided by a conditional variance.
@@ 4.1 Choice of Representation for Program Evaluation
This connection has a direct consequence for empirical practice: to improve efficiency with the RSV, we should not only predict the outcome $Y$ using observational data, but also predict the treatment $D$ using experimental data, and predict the sample indicator $S$ using all data. The first prediction is part of common practice, but the second and third are not. By using all three predictions, we use the RSV more efficiently.
@@
We confirm this connection by further developing the expression for $H^{*}(R)$: the numerator $\mathbb{E}(\Delta^{o}\mid R)$ contains $\Pr(Y=1,S=o|R)$ while the denominator $\sigma^{2}(\theta,R)$ contains $\Pr(D=1,S=e|R)$; see Lemma 2 below. Finally, by the definition of conditional probability, $\Pr(Y=1,S=o|R)=\Pr(Y=1|S=o,R)\Pr(S=o|R)$ and $\Pr(D=1,S=e|R)=\Pr(D=1|S=e,R)\Pr(S=e|R)$.
@@ 4.2 Inference with Learned Representations
For simplicity, we describe our inferential procedure using sample splitting with train and test folds, though our proofs in Appendix D allow for cross-fitting with any fixed number of folds. We first state our inferential procedure at a high level before filling in the details:
@@
    Learn the representation on train: $\widehat{H}(R)$.
@@
    • Construct an initial estimate, by regressing $\widehat{\mathbb{E}}(\Delta^{e}|R)$ on $\widehat{\mathbb{E}}(\Delta^{o}|R)$: $\widehat{\theta}_{\textsc{init}}$.
    • Combine these into the representation: $\widehat{H}(R)$.
@@
    Construct a causal estimate on test: $\widehat{\theta}$.
The overall structure is familiar (Angrist et al., 1999; Chernozhukov et al., 2018), although here we will show that no rate condition is required on $\widehat{H}(R)$. To simplify the algorithm statement, we introduce some notation for simple events and an algebraic identity.
Let $\mathbb{E}_{\textsc{train}}(\cdot)=\frac{1}{|\textsc{train}|}\sum_{i\in\textsc{train}}(\cdot)$ and $\mathbb{E}_{\textsc{test}}(\cdot)=\frac{1}{|\textsc{test}|}\sum_{i\in\textsc{test}}(\cdot)$. Slightly abusing notation, the counting operation, $\textsc{count}_{\textsc{event}}$, can be $\mathbb{E}_{\textsc{train}}(1_{\textsc{event}})$ in the second step or $\mathbb{E}_{\textsc{test}}(1_{\textsc{event}})$ in the third step, which will be clear from the context.
Lemma 2.
$(\Delta^{e}-\Delta^{o}\theta)^{2}=\frac{1\{D=1,S=e\}}{\Pr(D=1,S=e)^{2}}+\frac{1\{D=0,S=e\}}{\Pr(D=0,S=e)^{2}}+\theta^{2}\left[\frac{1\{Y=1,S=o\}}{\Pr(Y=1,S=o)^{2}}+\frac{1\{Y=0,S=o\}}{\Pr(Y=0,S=o)^{2}}\right].$
Algorithm 1 (Inference).
@@
    Learn the representation on train: $\widehat{H}(R)$.
@@
      Learn the representation: $\widehat{H}(R)=\frac{\widehat{\mathbb{E}}(\Delta^{o}|R)}{\widehat{\sigma}^{2}(\widehat{\theta}_{\textsc{init}},R)}$ where $\widehat{\sigma}^{2}(\widehat{\theta}_{\textsc{init}},R)$ is constructed from the marginal probabilities, predictors, and initial estimate via Lemma 2.
@@
    Construct a causal estimate on test: $\widehat{\theta}$.
@@
      Construct a causal estimate: $\widehat{\theta}=\frac{\mathbb{E}_{\textsc{test}}\{\widehat{\Delta}^{e}\widehat{H}(R)\}}{\mathbb{E}_{\textsc{test}}\{\widehat{\Delta}^{o}\widehat{H}(R)\}}$ where $\widehat{\Delta}^{e}$ and $\widehat{\Delta}^{o}$ are constructed from marginal probabilities according to (1).
@@
      Bootstrap its confidence interval: $\widehat{\theta}\pm c_{\alpha}\hat{v}n^{-1/2}$, where $c_{\alpha}$ is the $1-\alpha/2$ quantile of the standard Gaussian and $\hat{v}n^{-1/2}$ is the bootstrap standard error of $\widehat{\theta}$ while fixing $\widehat{H}(R)$.
@@
Sample splitting may be eliminated under complexity restrictions that tolerate simple machine learning procedures. See e.g., Chernozhukov et al. (2020) for a recent summary.
4.3 Robustness to Misspecification
In empirical research, the RSV is typically unstructured and high-dimensional, e.g., a satellite image. The prediction is often conducted by a complex machine learning algorithm, e.g., a deep convolutional neural network. For this realistic setting, rates of convergence are often unknown. In other words, we have no reason to believe that $\textsc{pred}_{Y}(R)$ converges to $\Pr(Y=1|S=o,R)$, nor that $\widehat{H}(R)$ converges to $H^{*}(R)$. Even carefully crafted architectures positing a generative model would typically be misspecified.
For this reason, we place a weaker regularity condition: the predictions, and hence the representation estimator, have some probability limit; they may be misspecified.
Assumption 4 (Limit).